A linear function has the same y-intercept as 4x + 5y = 20 and its graph contains the point (-1,-2). Find the y-intercept and slope of the linear function.
a. y-intercept: (0,4) slope: -6
b. y-intercept: (5,0) slope: -6
c. y-intercept: (0,4) slope: 6
d. y-intercept: (0,5) slope: 6

Answer by mananth(16946)


A linear function has the same y-intercept as 4x + 5y = 20
5y=-4x+20
y=-4/5 x +4
y intercept = 4
point (-1,-2)
y=mx+c
-2 = -1*m+4
-m=-6
m=6
c. y-intercept: (0,4) slope: 6